OpenStudy (anonymous):

Find the absolute max value and absolute min value of the function. (a) \[ g(x)=x^2+\frac{ 16}{ x } , 1\le x \le 4\] (b)\[f(x)=\left( x^2-9 \right)^{2/3}, -6\le x \le6\]

zepdrix (zepdrix):

\[\Large\rm g(x)=x^2+\frac{16}{x},\qquad x\in[1,4]\]We'll write g(x) like this,\[\Large\rm g(x)=x^2+16x^{-1},\qquad x\in[1,4]\]So again, I applied an exponent rule to make it look like that.

zepdrix (zepdrix):

This should make it easier to apply our power rule, yes?

OpenStudy (anonymous):

Yes

zepdrix (zepdrix):

So what do we get for g'(x)?

OpenStudy (anonymous):

g'(x)=2x-16x^-2

zepdrix (zepdrix):

\[\Large\rm g'(x)=2x-16x^{-2}\]\[\Large\rm g'(x)=2x-\frac{16}{x^2}\]Ok good. Let's get a common denominator,\[\Large\rm g'(x)=\frac{2x^3-16}{x^2}\]

zepdrix (zepdrix):

We'll look for critical points,\[\Large\rm 0=\frac{2x^3-16}{x^2}\]

zepdrix (zepdrix):

The denominator actually gives us some information, just not what we're looking for. It tells us where we can find `vertical tangents`, but we want `horizontal tangents` to the curve.

zepdrix (zepdrix):

So that leads us to,\[\Large\rm 0=2x^3-16\]

zepdrix (zepdrix):

Solve for x, what do you get? :o

OpenStudy (anonymous):

x=2!

zepdrix (zepdrix):

Ok good good good.

zepdrix (zepdrix):

Recall that when we're in a closed interval, we also need to check the endpoints!

zepdrix (zepdrix):

From here it's pretty simple: We'll plug the end points into the function, we'll plug the critical point into the function, largest output = our maximum smallest output = our minimum

zepdrix (zepdrix):

\[\Large\rm g(x)=x^2+\frac{16}{x}\] \[\Large\rm g(1)=?\]\[\Large\rm g(2)=?\]\[\Large\rm g(4)=?\]

OpenStudy (anonymous):

g(1)= 17 g(2)=12 (absolute min) g(4)=20 (absolute max)

zepdrix (zepdrix):

mmm ok good job! :)

OpenStudy (anonymous):

Need some help on the second one, please.

zepdrix (zepdrix):

\[\Large\rm f(x)=\left( x^2-9 \right)^{2/3}, \qquad x\in[-6,6]\]Were you able to find a derivative ok?

OpenStudy (anonymous):

\[f'(x)=2/3(x^2-9)^{-1/3}(2x)\] I'm not sure if this is correct?

zepdrix (zepdrix):

\[\Large\rm f'(x)=\frac{2}{3}(x^2-9)^{-1/3}(2x)\]Mmmm ok looks good.

zepdrix (zepdrix):

Let's condense it down a tad bit,\[\Large\rm f'(x)=\frac{4x}{3(x^2-9)^{1/3}}\]

zepdrix (zepdrix):

Oh a note about the last problem, because it's going to affect what we do here...

zepdrix (zepdrix):

I should not have thrown out the denominator so quickly in the last problem. We should have considered that point. It would have led to a different type of critical point at x=0. But since x=0 was NOT in our interval 1 to 4, we could easily throw it away. But not for the reason I mentioned before.

zepdrix (zepdrix):

So in this problem, we'll look for critical points as we did before,\[\Large\rm 0=\frac{4x}{3(x^2-9)^{1/3}}\]Let's just start by multiplying both sides by 3/4,\[\Large\rm 0=\frac{x}{(x^2-9)^{1/3}}\]

zepdrix (zepdrix):

We want to look at the denominator and numerator separately. Setting the denominator equal to 0 will give us these "other" types of critical points,\[\Large\rm 0=(x^2-9)^{1/3}\] What values of x do we get here?

OpenStudy (anonymous):

x=3?

zepdrix (zepdrix):

Mmm well you eventually get to,\[\Large\rm 0=x^2-9\]\[\Large\rm 9=x^2\]And to isolate x we need to square root, when we square root a square, we get the ole fancy plus/minus yes?

zepdrix (zepdrix):

The alternative is to apply your difference of squares rule from this step,\[\Large\rm 0=x^2-9\]\[\Large\rm 0=(x-3)(x+3)\]

OpenStudy (anonymous):

oh yeah! Forgot about that!

zepdrix (zepdrix):

So we get two values for x, x=-3, x=3. Now we'll set the numerator equal to zero and see what's going on there.

zepdrix (zepdrix):

That gives us a nice simple value x=0, yah?

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

So again, we'll plug our `end points` into the function, we'll plug our `critical points` into the function, and compare.

zepdrix (zepdrix):

\[\Large\rm f(x)=\left( x^2-9 \right)^{2/3}\] \[\Large\rm f(-6)=?\]\[\Large\rm f(-3)=?\]\[\Large\rm f(0)=?\]\[\Large\rm f(3)=?\]\[\Large\rm f(6)=?\]

zepdrix (zepdrix):

Here is what the function looks like graphed, https://www.desmos.com/calculator/rfuub7idg0 Spoiler Alert though. hehe

OpenStudy (anonymous):

f(-6)=9 f(-3)=0

OpenStudy (anonymous):

f(0)=4.3... f(3)=0

OpenStudy (anonymous):

f(6)=9

OpenStudy (anonymous):

Okay, thank you so much, zepdrix! :)

OpenStudy (anonymous):

I finally finished my math homework! YAY!!

zepdrix (zepdrix):

yay \c:/